If Ms. Whosier receives $16,000 in 4 years’ time, invests it in a bank account for 6 years at 6% per year, then 5 years at 3% per year, then 2 years at 4%, what sum will she have at the end of this 17 year period?
Answer #1

Ms. Whosier receives money in 4 years time, so actual investment is for 17-4 = 13 years only.

It is an illustration  of future value of single sum with different rates with different time period.

the formula is :

FV = PV * (1+r1)^n1 * (1+r2)^n2* (1+r3)^n3

FV = 16000 *( 1+0.06)^6 * (1+0.03)^5 * (1+0.04)^2

FV = 16000 * 1.418519 * 1.159274 * 1.0816 = 28458.23

Answer : 28458.23 (Thumbs up please)